The force R is the resultant of the forces P1, P2, and P3 acting on the rectangular plate. Find P1 and P2 if ๐‘… = 40 kN and ๐‘ƒ3 = 20 kN.


1
Expert's answer
2021-10-12T16:07:19-0400
"\vec P_1=\bigg(P_1\cos[\tan^{-1}(\dfrac{600}{300})], P_1\sin[\tan^{-1}(\dfrac{600}{300})]\bigg)"

"\vec P_2=\bigg(P_2\cos[\tan^{-1}(\dfrac{400}{300})], -P_2\sin[\tan^{-1}(\dfrac{400}{300})]\bigg)"

"\vec P_3=\bigg(-P_3, 0\bigg)"

"\vec R=\bigg(R\cos[30\degree], R\sin[30\degree]\bigg)"

"\vec R=\vec P_1+\vec P_2+\vec P_3"

"\cos[\tan^{-1}(\dfrac{600}{300})]=\cos[\tan^{-1}(2)]=\dfrac{1}{\sqrt{5}}"

"\sin[\tan^{-1}(\dfrac{600}{300})]=\sin[\tan^{-1}(2)]=\dfrac{2}{\sqrt{5}}"

"\cos[\tan^{-1}(\dfrac{400}{300})]=\cos[\tan^{-1}(\dfrac{4}{3})]=\dfrac{3}{5}"

"\sin[\tan^{-1}(\dfrac{400}{300})]=\sin[\tan^{-1}(\dfrac{4}{3})]=\dfrac{4}{5}"

"\cos[30\degree]=\dfrac{\sqrt{3}}{2}, \sin[30\degree]=\dfrac{1}{2}"

"40(\dfrac{\sqrt{3}}{2})=\dfrac{1}{\sqrt{5}}P_1+\dfrac{3}{5}P_2-20"

"40(\dfrac{1}{2})=\dfrac{2}{\sqrt{5}}P_1-\dfrac{4}{5}P_2+0"



"\dfrac{1}{\sqrt{5}}P_1=20\sqrt{3}+20-\dfrac{3}{5}P_2"

"\dfrac{1}{\sqrt{5}}P_1=10+\dfrac{2}{5}P_2"

"20\sqrt{3}+20-\dfrac{3}{5}P_2=10+\dfrac{2}{5}P_2"

"P_2=20\sqrt{3}+30"

"\dfrac{1}{\sqrt{5}}P_1=10+\dfrac{2}{5}(20\sqrt{3}+30)"

"P_1=8\sqrt{15}+12\sqrt{5}"

"P_1=4\sqrt{5}(2\sqrt{3}+3) \ kN"

"P_2=10(2\sqrt{3}+3)\ kN"